# Ov3R: Open-Vocabulary Semantic 3D Reconstruction from RGB Videos

### v1

@@ Abstract
We present Ov3R, a novel framework for open-vocabulary semantic 3D reconstruction from RGB video streams, designed to advance Spatial AI. The system features two key components: CLIP3R, a CLIP-informed 3D reconstruction module that predicts dense point maps from overlapping clips while embedding object-level semantics; and 2D–3D OVS, a 2D-3D open-vocabulary semantic module that lifts 2D features into 3D by learning fused descriptors integrating spatial, geometric, and semantic cues. Unlike prior methods, Ov3R incorporates CLIP semantics directly into the reconstruction process, enabling globally consistent geometry and fine-grained semantic alignment. Our framework achieves state-of-the-art performance in both dense 3D reconstruction and open-vocabulary 3D segmentation — marking a step forward toward real-time, semantics-aware Spatial AI.
@@ 1 Introduction
Spatial AI systems (Davison 2018) aim to understand both the geometry and semantics of the surrounding environment from images in real-time, enabling an embedded AI agent to navigate and interact with it effectively.
@@
Dense 3D reconstruction lies at the core of these systems. Over the decades, this task has taken various forms — from Structure-from-Motion (SFM) (Lindenberger et al. 2021; Liu et al. 2023; Schonberger and Frahm 2016; Snavely, Seitz, and Szeliski 2006) and Multi-View Stereo (MVS), typically performed offline, to Simultaneous Localization and Mapping (SLAM) (Campos et al. 2021; Engel, Schöps, and Cremers 2014; Gong et al. 2025; Teed and Deng 2021), which is executed online by autonomous agents and thus requires real-time processing. The latter flavor is the most suitable approach for developing Spatial AI systems, although it poses greater challenges compared to offline methods, as input images are collected incrementally rather than being available in advance - making the system more susceptible to noise, motion blur, and cumulative drift over time.
@@
In the last few years, SLAM has undergone a resurgence, fueled by advancements in related fields such as novel view synthesis (Mildenhall et al. 2021; Kerbl et al. 2023) and, more recently, the emergence of 3D reconstruction models (Wang et al. 2024b) (3R), which have redefined the classical reconstruction paradigm by bypassing explicit camera pose estimation. The former has led to a family of SLAM methods, based on NeRF or 3DGS, yielding dense 3D reconstructions and accurate camera tracking, albeit with high computational cost due to the need for scene-specific training. The latter, in contrast, has resulted in the first 3R-based systems capable of real-time performance, though often at the expense of tracking accuracy, due to the absence of explicit pose estimation. In both directions, however, the semantic understanding of the surrounding environment has remained largely unexplored – and where considered, it has been limited to simplified scenarios with a predefined set of classes (Zhu et al. 2024c, b). As a result, a significant gap between existing SLAM methods and envisioned Spatial AI systems still persists.
@@
In parallel, with the emergence of CLIP (Radford et al. 2021), research on open-vocabulary 3D semantic understanding has surged. However, existing approaches largely rely on offline reconstruction pipelines (Kerr et al. 2023; Nguyen et al. 2024; Peng et al. 2023; Qin et al. 2024; Takmaz et al. 2023; Tie et al. 2024) or RGBD SLAM methods that require depth sensors (Martins, Oswald, and Civera 2024), and therefore do not address the aforementioned gap.
@@
In this paper, we introduce Ov3R, an open-vocabulary semantic 3D reconstruction framework that processes RGB-only video streams. Ov3R comprises two key components to achieve both accurate 3D reconstruction and semantic scene understanding: (i) CLIP3R, a CLIP-informed 3R model enriched with object-level CLIP features that encode the semantics of individual objects in the scene; and (ii) 2D-3D OVS, a module that integrates rich semantic information from various foundation models — CLIP, DINO and a distilled 3D-DINO encoder — to effectively segment the dense 3D reconstruction produced by CLIP3R. These two components are loosely coupled through the CLIP features, making Ov3R a modular framework capable of performing the two tasks independently – e.g., running only 3D reconstruction when semantic understanding is not required.
@@
  • Ov3R achieves state-of-the-art performance in both 3D reconstruction and open-vocabulary 3D segmentation tasks.
@@ 2 Related Work
Offline Pipelines. This category includes strategies that assume the availability of a complete set of images beforehand (Lindenberger et al. 2021; Liu et al. 2023; Schonberger and Frahm 2016; Snavely, Seitz, and Szeliski 2006), from which to process either camera parameters and 3D points via Structure from Motion (SfM) (Arrigoni 2025), or 3D points only through Multi-View Stereo (MVS) (Wang et al. 2024a). With the advent of NeRF (Mildenhall et al. 2021) and 3DGS (Kerbl et al. 2023) for novel view synthesis, several derived frameworks (Li et al. 2023b; Chen et al. 2024b, c; Guédon and Lepetit 2024; Chen et al. 2024a) have been developed to achieve high-fidelity 3D reconstructions. However, offline processing does not fulfill the requirements of a Spatial AI system.
SLAM Systems. These approaches estimate the camera pose and the 3D structure of the environment during navigation – i.e., while images are being collected in real-time. Early hand-crafted works (Campos et al. 2021; Teed and Deng 2021; Bailey and Durrant-Whyte 2006; Durrant-Whyte and Bailey 2006; Engel, Schöps, and Cremers 2014) focused mostly on camera tracking, while yielding sparse 3D reconstructions. More recently, the integration of novel views synthesis techniques (NeRF and 3DGS) into SLAM (Tosi et al. 2024) has enabled both accurate and dense reconstructions, either assuming the availability of depth sensors (Zhang et al. 2023; Li, Liu, and Wu 2024; Li et al. 2021) or not (Zhang et al. 2023; Zhu et al. 2024d; Li et al. 2023a; Zhu et al. 2024a; Li et al. 2023a). Although some of these frameworks have been enhanced with semantics understanding (Zhu et al. 2024c, b), their scope is often limited to a fixed set of classes and lacks real-time capabilities.
3R-driven Models. DUSt3R (Wang et al. 2024b) pioneered the first end-to-end model for 3D reconstruction from two uncalibrated RGB images, predicting two pointmaps aligned to the same reference system, from which intrinsic and extrinsic camera parameters can also be derived. Several extensions of DUSt3R target explicit features matching (Leroy, Cabon, and Revaud 2024), novel view synthesis (Wang et al. 2025), dynamic scene reconstructions (Zhang et al. 2024), although not achieving real-time performance. To address these limitations, Spann3R (Wang and Agapito 2024) introduces an incremental reconstruction approach without global alignment, enabling fast real-time operation. However, it rapidly accumulates drift errors and produces poor reconstruction accuracy. On the same track, SLAM3R (Liu et al. 2025) revisits SLAM by proposing a real-time 3D reconstruction method that does not require explicit pose estimation. Nonetheless, current 3R-based frameworks remain focused on detailed geometry reconstruction, disregarding higher-level scene understanding.
@@
2 3D Open Vocabulary Semantics
Closed-vocabulary Online Semantics. These methods predict per-pixel segmentation using pre-defined classes. SemanticFusion pioneered per-pixel 3D semantic prediction from multi-view images. Recently, several semantic SLAM methods have been introduced to simultaneously reconstruct dense maps and semantics. NIS-SLAM (Zhai et al. 2024) integrates semantics into tri-plane features to optimize both scene representation and semantic predictions. SGS-SLAM (Li et al. 2024) and NEDS-SLAM (Ji et al. 2024) encode semantics into Gaussians and incrementally reconstruct scenes with a closed-set segmentation. These methods achieve scene understanding by assuming a known set of 2D semantic annotations. However, relying on pre-defined semantic classes limits their applicability in real-world scenarios where novel object categories may appear.
3D Open-Vocabulary Offline Semantics. Most open-vocabulary semantic methods rely heavily on pre-processed 3D scene points as input. OpenScene (Peng et al. 2023) encodes per-pixel dense features in CLIP feature space and associates 2D dense features with known 3D scene points. Open3DIS (Nguyen et al. 2024) relies significantly on pre-processed 2D and 3D semantic instances to extract pointwise features. Unlike these approaches that require known 3D points, OpenNeRF (Engelmann et al. 2024) directly uses RGB images as input for NeRF and OpenSeg (Ghiasi et al. 2022). Similarly, HOV-SG (Werby et al. 2024) extracts 3D semantics and scene graph hierarchy from RGB-D sequences, relying heavily on feature pre-processing. Unlike the above offline methods, OVO (Martins, Oswald, and Civera 2024) proposes the first online 3D open-vocabulary semantic mapping compatible with SLAM systems. It can extract and track 3D segments from RGB-D videos while encoding CLIP descriptors through a proposed CLIP merging strategy. However, it still requires depth information and explicit pose estimation from SLAM systems. Besides, OVO computes CLIP descriptors solely from 2D images, overlooking the potential of 3D geometric information. In contrast, our Ov3R can extract 3D segments without depth sensors or explicit camera tracking. Additionally, we introduce a 2D-3D OVS module integrating DINO features and 3D information into CLIP descriptors.
@@ 3 Method
We now introduce Ov3R, whose overall architecture is depicted in Figure 2. It consists of two main components, highlighted in red in the figure: (i) a CLIP-informed 3R-based model (CLIP3R) and (ii) a 2D-3D OVS module. CLIP3R produces dense scene points with semantic alignment in real-time, while 2D-3D OVS serves as a downstream task – open-vocabulary semantic segmentation. It enables Ov3R to reconstruct 3D semantic scenes while retaining the possibility to perform two tasks independently.
1 CLIP3R
@@
$\Phi_{3R}\left(I_{i}^{(H\times W\times 3)}\right)\to P_{i}^{(H\times W\times 3)},i=1,...,L$ (1)
@@
As shown in Figure 3, inspired by (Liu et al. 2025), CLIP3R deploys two main branches: an Image-to-Points (I2P, in yellow) model and a Local-to-World (L2W, in blue) model.
@@
1) The former is a ViT, inspired by DUSt3R that runs over a local window of images to predict a pointmap aligned to the reference system of the central frame (i.e., the keyframe). This is achieved by processing images with a shared encoder $E_{img}$ and two decoders $D_{key}$ and $D_{sup}$ for the keyframe and remaining images, respectively. Notably, no explicit camera pose estimation is performed, which in turn can be derived from the pointmap as a byproduct.
@@
2) The latter, instead, processes the local pointmaps and aligns them to the scene-level reconstruction. It employs a pointmap encoder $E_{pts}$, a registration decoder $D_{reg}$, and a scene decoder $D_{sce}$, sharing the same structure as the encoder and decoder used by I2P. It identifies correlated keyframes using a reservoir and retrieval strategy (Liu et al. 2025) and processes them to predict global scene points $\hat{P^{\prime}_{i}}$.
@@
Both branches, however, lack high-level reasoning about scene semantics, which we address through integration of CLIP features and CLIP-semantic supervision, to learn a stronger semantic-geometric relationship and semantic consistency.
@@
Object-level CLIP features. Although rich in semantics, CLIP embeddings are image-level features and therefore fail at modeling the fine-grained semantics of individual objects in the scene (Zuo et al. 2025). To overcome this limitation, we extract object-level CLIP features $f_{\text{oCLIP}}^{\left(H\times W\times D\right)}$ in place of vanilla CLIP features. This is achieved by first obtaining $M$ object masks $m^{\left(H\times W\times 1\right)}$ using the Segment Anything Model (SAM) (Kirillov et al. 2023). We then create $M$ masked images and process them to extract CLIP patch embeddings, which are averaged and upsampled to $f_{\text{CLIP}}^{\left(H\times W\times D\right)}$ to match the image resolution. Finally, the object-level features $f_{\text{oCLIP}}^{\left(H\times W\times D\right)}$ are obtained by combining the individual CLIP features obtained from the $M$ masked images within a single features map as:
@@
$f_{\text{oCLIP}}^{\left(H\times W\times D\right)}=\sum_{i=0}^{M}m_{i}^{\left(H\times W\times 1\right)}\cdot f_{\text{CLIP}_{i}}^{\left(H\times W\times D\right)}$ (2)
@@
These finer-level features are then integrated into CLIP3R.
@@
CLIP-informed I2P. To predict local pointmaps, the I2P branch extracts visual tokens $F_{vit}^{\left(T\times d\right)}$ from each frame in the processed window through ViT blocks (Dosovitskiy et al. 2020). To enrich I2P with knowledge from CLIP, we concurrently extract object-level CLIP features and tokenize them into $F_{oCLIP}^{\left(T\times d\right)}$. Then, $F_{oCLIP}$ and $F_{vit}$ are processed through cross-attention to obtain the fused features $F_{fuse}$, which are added to $F_{vit}$. The encoding process is denoted as:
@@
$F_{fuse}=F_{vit}+softmax(\frac{F_{vit}F_{oCLIP}}{\sqrt{F_{oCLIP}}})F_{oCLIP}$ (3)
@@
These features are then processed by the keyframe decoder $D_{key}$ and the supporting decoder $D_{sup}$ from the original I2P. A linear head is applied to predict both scene points $P^{\prime}_{i}$ ($P^{\prime}_{k},P^{\prime}_{s}\in P^{\prime}_{i}$) and confidence map $C$. The revised I2P network is trained end-to-end through a confidence-aware loss over ground truth scene points:
$L_{I2P}=\sum_{i=1}^{L}M_{i}\cdot(C\cdot L1(\frac{1}{z^{\prime}}P^{\prime}_{i},\frac{1}{z}P_{i})-\alpha\log C)$ (4)
@@
where $L$ is the number of supporting frames, $M_{i}$ are the masked points with ground truth values in $P_{i}$, $L1$ is the point-wise Euclidean distance, $z$ and $z^{\prime}$ are scale factors, and $\alpha$ is a hyperparameter controlling the regularization term.
Figure 3: CLIP3R Overview. I2P integrates object-level CLIP features with visual embeddings to predict local pointmaps. L2W then aligns these local pointmaps to global scene coordinates while predicting object-level CLIP features in scene space.
CLIP-informed L2W. The L2W branch maintains its original structure but incorporates an additional prediction head for object-level CLIP features $f_{oCLIP}^{{}^{\prime}}$. This head embeds fine-grained semantic knowledge directly into the final 3D reconstruction while enforcing semantic consistency across the scene. This head is implemented as a DPT (Ranftl, Bochkovskiy, and Koltun 2021), and the entire L2W branch is trained using two losses: $L_{L2W}$ and $L_{oCLIP}$. The former is similar to the loss used to supervise I2P:
$L_{L2W}=\sum_{i=1}^{L}M_{i}\cdot(\hat{C}\cdot L1(\hat{P^{\prime}_{i}},P_{i})-\alpha\log\hat{C})$ (5)
while the latter minimizes the feature distance between extracted $f_{oCLIP}$ and predicted $f_{oCLIP}^{{}^{\prime}}$ features:
$L_{oCLIP}=\left\|f_{oCLIP}^{{}^{\prime}}-f_{oCLIP}\right\|_{1}$ (6)
@@
Figure 4: 2D-3D OVS Overview. After matching 2D and 3D segments across images and pointmaps, CLIP, DINO, and 3D features are combined into a 2D-3D fused descriptor.
@@
2 2D-3D OVS
3D Segments Matching. We adopt the strategy from (Martins, Oswald, and Civera 2024): after predicting dense point maps for each frame, we generate 2D object masks from images using SAM (Kirillov et al. 2023). We then project the 3D point map onto each keyframe and assign a 2D segment labels, corresponding to the SAM-predicted masks, to the matched 3D points. After processing all 2D segments, 3D points sharing the same label are merged into coherent 3D segments. However, this strategy relies solely on 2D information and lacks explicit semantic knowledge from the 3D geometry itself. To address this limitation, we introduce 2D-3D fused descriptors, obtained as follows.
2D-3D Fused Descriptors. CLIP features alone are insufficient to model semantics in 3D space, as they lack fine-grained details and clear object boundaries. While DINO (Caron et al. 2021) features extract clear boundaries between inter-components, they lack knowledge about 3D geometry – which is, on the contrary, available from point features (Hegde, Valanarasu, and Patel 2023) extracted from 3D pointclouds. Therefore, we introduce a 2D-3D fused descriptor that combines these three complementary features types extracted from CLIP, DINO, and a 3D-CLIP encoder (Hegde, Valanarasu, and Patel 2023).
To capture contextual information, we adopt the multi-level architecture by (Werby et al. 2024) and (Martins, Oswald, and Civera 2024), as shown in Figure 4. From each keyframe, we extract CLIP and DINO features at three levels: from the full frame $I_{full}$ and from a crop containing the 2D segment (w/ background) $I_{seg}$, while encoding CLIP and 3D points features after masking the crop to remove background $I_{oseg}$ (w/o background), and its corresponding portion of the pointmap $P_{oseg}$. In the first two branches, the DINO features $F_{dino}$ ($F_{dino}^{f},F_{dino}^{s}\in F_{dino}$) extracted though ViT-small architecture (Caron et al. 2021) are first projected to the same feature dimension as CLIP features by a linear layer. The DINO features are then concatenated with CLIP features $F_{clip}$ ($F_{clip}^{f},F_{clip}^{s}\in F_{clip}$), followed by a linear layer to transform the dimensions:
$F_{cat}=Linear([F_{dino},F_{clip}])$ (7)

### v2

@@ Abstract
We present Ov3R, a novel framework for open-vocabulary semantic 3D reconstruction from RGB video streams, designed to advance Spatial AI. The system features two key components: CLIP3R, a CLIP-informed 3D reconstruction module that predicts dense point maps from overlapping clips alongside object-level semantics; and 2D–3D OVS, a 2D-3D open-vocabulary semantic module that lifts 2D features into 3D by learning fused descriptors integrating spatial, geometric, and semantic cues. Unlike prior methods, Ov3R incorporates CLIP semantics directly into the reconstruction process, enabling globally consistent geometry and fine-grained semantic alignment. Our framework achieves state-of-the-art performance in both dense 3D reconstruction and open-vocabulary 3D segmentation — marking a step forward toward real-time, semantics-aware Spatial AI.
@@ 1 Introduction
Spatial AI systems [davison2018futuremapping] aim to understand both the geometry and semantics of the surrounding environment from images in real-time, enabling an embedded AI agent to navigate and interact with it effectively.
@@
Dense 3D reconstruction lies at the core of these systems. Over the decades, this task has taken various forms — from Structure-from-Motion (SFM) [lindenberger2021pixel, liu20233d, schonberger2016structure, snavely2006photo] and Multi-View Stereo (MVS), typically performed offline, to Simultaneous Localization and Mapping (SLAM) [campos2021orb, engel2014lsd, gong2025hs, teed2021droid], which is executed online by autonomous agents and thus requires real-time processing. The latter flavor is the most suitable approach for developing Spatial AI systems, although it poses greater challenges compared to offline methods, as input images are collected incrementally rather than being available in advance – making the system susceptible to noise, motion blur, and cumulative drift over time.
@@
In the last few years, SLAM has undergone a resurgence, fueled by advancements in related fields such as novel view synthesis [mildenhall2021nerf, kerbl20233d] and, more recently, the emergence of 3D reconstruction models [wang2024dust3r] (3R), which have redefined the classical reconstruction paradigm by bypassing explicit camera pose estimation. The former has led to a family of SLAM methods, based on NeRF or 3DGS, yielding dense 3D reconstructions and accurate camera tracking, albeit with high computational cost due to the need for scene-specific training. The latter, in contrast, has resulted in the first 3R-based systems capable of real-time performance, though often at the expense of tracking accuracy, due to the absence of explicit pose estimation. In both directions, however, the semantic understanding of the surrounding environment has remained largely unexplored – and where considered, it has been limited to simplified scenarios with a predefined set of classes [zhu2024sni, zhu2024semgauss]. As a result, a significant gap between existing SLAM methods and envisioned Spatial AI systems still persists.
@@
In parallel, with the emergence of CLIP [radford2021learning], research on open-vocabulary 3D semantic understanding has surged. However, existing approaches largely rely on offline reconstruction pipelines [kerr2023lerf, nguyen2024open3dis, peng2023openscene, qin2024langsplat, takmaz2023openmask3d, tie20242] or RGBD SLAM methods that require depth sensors [martins2024ovo], and therefore do not address the aforementioned gap.
@@
In this paper, we introduce Ov3R, an open-vocabulary semantic 3D reconstruction framework that processes RGB-only video streams. Ov3R comprises two key components to achieve both accurate 3D reconstruction and semantic scene understanding: (i) CLIP3R, a CLIP-informed 3R model enriched with object-level CLIP features that encode the semantics of individual objects in the scene; and (ii) 2D-3D OVS, a module that integrates rich semantic information from various foundation models — CLIP3R itself, DINO and a distilled 3D-DINO encoder — to effectively segment the dense 3D reconstruction produced by CLIP3R. These two components are loosely coupled through the CLIP features, making Ov3R a modular framework capable of performing the two tasks either jointly or independently – e.g., running only 3D reconstruction when semantic understanding is not required or vice-versa.
@@
  • Ov3R achieves state-of-the-art performance in 3D reconstruction and open-vocabulary 3D segmentation.
@@ 2 Related Work
Offline Pipelines. This category includes strategies that assume the availability of a complete set of images beforehand [lindenberger2021pixel, liu20233d, schonberger2016structure, snavely2006photo], from which to process either camera parameters and 3D points via Structure from Motion (SfM) [arrigoni2025taxonomy], or 3D points only through Multi-View Stereo (MVS) [wang2024learning]. With the advent of NeRF [mildenhall2021nerf] and 3DGS [kerbl20233d] for novel view synthesis, several derived frameworks [li2023neuralangelo, chen2024mvsplat, chen2024mvsplat360, guedon2024sugar, chen2024pgsr] have been developed to achieve high-fidelity 3D reconstructions. However, offline processing does not fulfill the requirements of a Spatial AI system.
SLAM Systems. These approaches estimate the camera pose and the 3D structure of the environment during navigation – i.e., while images are being collected in real-time. Early hand-crafted works [campos2021orb, teed2021droid, bailey2006simultaneous, durrant2006simultaneous, engel2014lsd] focused mostly on camera tracking, while yielding sparse 3D reconstructions. More recently, the integration of novel views synthesis techniques (NeRF and 3DGS) into SLAM [tosi2024nerfs] has enabled both accurate and dense reconstructions, either assuming the availability of depth sensors [zhang2023go, li2024cto, li2021po] or not [zhang2023go, zhu2024nicer, li2023dense, zhu2024mgs, li2023dense]. Although some of these frameworks have been enhanced with semantics understanding [zhu2024sni, zhu2024semgauss], their scope is often limited to a fixed set of classes and lacks real-time capabilities.
3R-driven Models. DUSt3R [wang2024dust3r] pioneered the first end-to-end model for 3D reconstruction from two uncalibrated RGB images, predicting two pointmaps aligned to the same reference system, from which intrinsic and extrinsic camera parameters can also be derived. Several extensions of DUSt3R target explicit features matching [leroy2024grounding], single-image metric depth estimation [wang2025moge], and dynamic scene reconstructions [zhang2024monst3r], although not achieving real-time performance. To address these limitations, Spann3R [wang20243d] introduces an incremental reconstruction approach to enable fast real-time operation. However, it rapidly accumulates drift errors due to the lack of global alignment and yields poor reconstruction accuracy. On the same track, SLAM3R [liu2025slam3r] revisits SLAM by proposing a real-time 3D reconstruction method that does not require explicit pose estimation. Nonetheless, current 3R-based frameworks remain focused on detailed geometry reconstruction, disregarding higher-level scene understanding.
@@
2.2 3D Semantics
We can broadly classify 3D semantic segmentation approaches into closed-vocabulary and open-vocabulary ones.
Closed-Vocabulary Online Semantics. These methods predict per-pixel segmentation using pre-defined classes. SemanticFusion [mccormac2017semanticfusion] pioneered per-pixel 3D semantic prediction from multi-view images. Recently, several semantic SLAM methods have been introduced to simultaneously reconstruct dense maps and semantics. NIS-SLAM [zhai2024nis] integrates semantics into tri-plane features to optimize both scene representation and semantic predictions. SGS-SLAM [li2024sgs] and NEDS-SLAM [ji2024neds] encode semantics into Gaussians and incrementally reconstruct scenes with a closed-set segmentation. These methods achieve scene understanding by assuming a known set of 2D semantic annotations. However, relying on pre-defined semantic classes limits their applicability in real-world scenarios where novel object categories may appear.
3D Open-Vocabulary Offline Semantics. Most open-vocabulary semantic methods rely heavily on pre-processed 3D scene points as input. OpenScene [peng2023openscene] encodes per-pixel dense features in CLIP feature space and associates 2D dense features with known 3D scene points. Open3DIS [nguyen2024open3dis] relies significantly on pre-processed 2D and 3D semantic instances to extract pointwise features. Unlike these approaches that require known 3D points, OpenNeRF [engelmann2024opennerf] directly uses RGB images as input for NeRF and OpenSeg [ghiasi2022scaling]. Similarly, HOV-SG [werby2024hierarchical] extracts 3D semantics and scene graph hierarchy from RGB-D sequences, relying heavily on feature pre-processing. Unlike the above offline methods, OVO [martins2024ovo] proposes the first online 3D open-vocabulary semantic mapping compatible with SLAM systems. It can extract and track 3D segments from RGB-D videos while encoding CLIP descriptors through a proposed CLIP merging strategy. However, it still requires depth information to be collected by an active sensor and explicit pose estimation from a SLAM system running concurrently. Besides, OVO computes CLIP descriptors solely from 2D images, overlooking the potential of 3D geometric information. In contrast, our Ov3R can extract 3D segments from RGB-only videos, not requiring explicit camera tracking.
@@ 3 Method
Figure 3: CLIP3R Overview. I2P integrates object-level CLIP features with visual embeddings to predict local pointmaps. L2W then aligns these local pointmaps to global scene coordinates while predicting object-level CLIP3R features with a DPT head in scene space.
We now introduce Ov3R, whose overall architecture is depicted in Figure 2. It consists of two main components, highlighted in yellow and blue: (i) a CLIP-informed 3R-based model (CLIP3R) and (ii) a 2D-3D OVS module. CLIP3R produces dense scene points with semantic alignment in real-time, while 2D-3D OVS processes the semantic features predicted by the former for the downstream task of open-vocabulary 3D semantic segmentation. It enables Ov3R to reconstruct 3D semantic scenes while retaining the possibility to perform two tasks either jointly or independently.
3.1 CLIP3R
@@
$\Phi_{3R}\left(I_{i}^{(N\times H\times W\times 3)}\right)\to P_{i}^{(H\times W\times 3)},i=1,...,L$ (1)
@@
As shown in Figure 3, inspired by [liu2025slam3r], CLIP3R deploys two main branches: an Image-to-Points (I2P, in yellow) model and a Local-to-World (L2W, in green) model.
@@
1) The former is a ViT, inspired by DUSt3R, running over a local window of images to predict a pointmap aligned to the reference system of the central frame (i.e., the keyframe). This is achieved by processing images with a shared encoder $E_{img}$ and two decoders $D_{key}$ and $D_{sup}$ for the keyframe and remaining images, respectively. Notably, no explicit camera pose estimation is performed, which in turn can be derived from the pointmap as a byproduct.
@@
2) The latter, instead, processes the local pointmaps and aligns them to the scene-level reconstruction. It employs a pointmap encoder $E_{pts}$, a registration decoder $D_{reg}$, and a scene decoder $D_{sce}$, sharing the same structure as the encoder and decoder used by I2P. It identifies correlated keyframes using a reservoir and retrieval strategy [liu2025slam3r] and processes them to predict global scene points $\hat{P^{\prime}_{i}}$.
@@
Both branches, however, lack high-level reasoning about scene semantics, which we address through integration of CLIP features and CLIP-semantic supervision, to learn stronger semantic-geometric relationships and consistency.
@@
Object-level CLIP3R features. Although rich in semantics, CLIP embeddings are image-level features and therefore fail at modeling the fine-grained semantics of individual objects in the scene [zuo2025fmgs]. To overcome this limitation, we extract object-level CLIP3R features $\mathbf{F}_{\text{oCLIP}}^{\left(H\times W\times D\right)}$ in place of vanilla CLIP features. This is achieved by first obtaining $M$ object masks $m^{\left(H\times W\times 1\right)}$ using the Segment Anything Model (SAM) [kirillov2023segment]. We then create $M$ masked images and process them to extract CLIP patch embeddings, which are averaged and upsampled to $\mathbf{F}_{\text{CLIP}}^{\left(H\times W\times D\right)}$ to match the image resolution. Finally, the object-level features $\mathbf{F}_{\text{oCLIP}}^{\left(H\times W\times D\right)}$ are obtained by combining the individual CLIP features obtained from the $M$ masked images within a single features map as:
@@
$\mathbf{F}_{\text{oCLIP}}^{\left(H\times W\times D\right)}=\sum_{i=0}^{M}m_{i}^{\left(H\times W\times 1\right)}\cdot\mathbf{F}_{\text{CLIP}_{i}}^{\left(H\times W\times D\right)}$ (2)
@@
These finer-level features are then integrated into CLIP3R, both in I2P and L2W branches in different ways.
@@
CLIP-informed I2P. To predict local pointmaps, the I2P branch extracts visual tokens $\mathbf{F}_{\text{ViT}}^{\left(T\times d\right)}$ from each frame in the processed window through ViT blocks [dosovitskiy2020image]. To enrich I2P with knowledge from CLIP, we concurrently extract object-level CLIP3R features and tokenize them into $\mathbf{F}_{\text{oCLIP}}^{\left(T\times d\right)}$. Then, $\mathbf{F}_{\text{oCLIP}}$ and $\mathbf{F}_{vit}$ are processed through cross-attention to obtain the fused features $\mathbf{F}_{\text{fuse}}$, which are added to $\mathbf{F}_{\text{ViT}}$. The encoding process is denoted as:
@@
$\mathbf{F}_{\text{fuse}}=\mathbf{F}_{\text{ViT}}+\texttt{softmax}(\frac{\mathbf{F}_{\text{ViT}}\cdot\mathbf{F}_{\text{oCLIP}}^{T}}{\sqrt{d}})\cdot\mathbf{F}_{\text{oCLIP}}$ (3)
@@
where $d$ denotes the feature dimension. These features are then processed by the keyframe decoder $D_{key}$ and the supporting decoder $D_{sup}$ from the original I2P. A linear head is applied to predict both scene points $P^{\prime}_{i}$ ($P^{\prime}_{k},P^{\prime}_{s}\in P^{\prime}_{i}$) and confidence map $C$. The revised I2P network is trained end-to-end through a confidence-aware loss over ground truth scene points:
[…]
@@
where $L$ is the number of supporting frames, $M_{i}$ are the masked points with ground truth values in $P_{i}$, $z$ and $z^{\prime}$ are scale factors, and $\alpha$ controls the confidence-based regularization term.
@@
Figure 4: 2D-3D OVS Overview. After matching 2D and 3D segments across images and pointmaps, CLIP3R, DINO, and 3D-CLIP features are combined into a 2D-3D fused descriptor, on top of which open-vocabulary semantic segmentation is performed. CLIP3R and DINO features are processed both at scene and instance levels. Meanwhile, 3D-CLIP features are extracted from masked 3D object points.
CLIP-informed L2W. The L2W branch maintains its original structure but incorporates an additional prediction head for object-level CLIP3R features, we refer to as $\mathbf{F}_{\text{CLIP3R}}$. This head embeds fine-grained semantic knowledge directly into the final 3D reconstruction while enforcing semantic consistency across the scene. This head is implemented as a DPT [ranftl2021vision], and the entire L2W branch is trained using two losses: $\mathcal{L}_{L2W}$ and $\mathcal{L}_{oCLIP}$. The former is similar to the loss used to supervise I2P:
[…]
while the latter minimizes the feature distance between extracted $\mathbf{F}_{\text{oCLIP}}$ and predicted $\mathbf{F}_{\text{CLIP3R}}$ features:
$\mathcal{L}_{\text{oCLIP}}=\left\|\mathbf{F}_{\text{CLIP3R}}-\mathbf{F}_{\text{oCLIP}}\right\|_{1}$ (6)
3.2 2D-3D OVS
3D Segments Matching. We revise and improve the strategy from [martins2024ovo], replacing CLIP features with those predicted by CLIP3R: after predicting dense point maps for each frame, we project the 3D point map predicted by CLIP3R onto each keyframe to assign 2D segment labels to the matched 3D points, corresponding to the SAM-predicted masks already used by CLIP3R itself. After processing all 2D segments, 3D points sharing the same label are merged into coherent 3D segments. Although this strategy effectively selects individual semantic components in the scene, it relies solely on 2D information and lacks explicit semantic knowledge from the 3D geometry itself. To address this limitation, we introduce 2D-3D fused descriptors, obtained as follows.
2D-3D Fused Descriptors. CLIP3R features alone are insufficient to model semantics in 3D space, as they are learned from 2D CLIP features and thus lack fine-grained 3D structural information. On the same track, while DINO [caron2021emerging] features extract clear boundaries between inter-components, these lack knowledge about 3D geometry – which is, on the contrary, available from point features [hegde2023clip] extracted from 3D pointclouds. Therefore, we introduce a 2D-3D fused descriptor that combines these three complementary feature types extracted from i) CLIP3R, ii) DINO, and iii) a 3D-CLIP encoder [hegde2023clip]:
i) To capture contextual information, we adopt the multi-level architecture by [werby2024hierarchical] and [martins2024ovo], as shown in Figure 4. From each keyframe, we leverage features predicted by CLIP3R (middle branch in the figure) at both scene-level (blue arrows in figure) and object instance-level (yellow arrows in figure), the latter by cropping the former according to 2D masks by SAM – referred to as $\mathbf{F}_{\text{CLIP3R}}^{\text{scene}}$ and $\mathbf{F}_{\text{CLIP3R}}^{\text{inst}}$;
ii) In parallel, DINO features are extracted (top branch in the figure) and again processed at both the scene and instance levels, similarly to CLIP3R features. These are referred to, respectively, as $\mathbf{F}_{\text{DINO}}^{\text{scene}}$ and $\mathbf{F}_{\text{DINO}}^{\text{inst}}$;
iii) Finally, instance-level 3D features are extracted from the point map after retaining only 3D points associated with the same object label (bottom branch in the figure). A 2D counterpart of the object-level features is retrieved from CLIP3R, by masking the cropped instance-level features according to 2D segmentation masks corresponding to the same object label (red arrow).
@@
Therefore, 2D-aware features by DINO and CLIP3R are concatenated at both the scene and the instance levels independently (© in the figure), after projecting the former to the same feature dimension as the latter by a linear layer. Concatenated features are further processed by a linear layer
$\mathbf{F}_{cat}^{\text{scene}}=\texttt{Linear}([\texttt{Linear}(\mathbf{F}_{\text{DINO}}^{\text{scene}}),\mathbf{F}_{\text{CLIP3R}}^{\text{scene}}])$ (7)
$\mathbf{F}_{cat}^{\text{inst}}=\texttt{Linear}([\texttt{Linear}(\mathbf{F}_{\text{DINO}}^{\text{inst}}),\mathbf{F}_{\text{CLIP3R}}^{\text{inst}}])$ (8)